OpenStudy (anonymous):

How do I find five ordered pairs for equation y=3x-2

OpenStudy (anonymous):

pick a number, any number i pick 2, then \(y=3\times 2-2=6-2=4\) so one ordered pair is \((2,4)\) lather, rinse, repeat